OpenStudy (anonymous):

CONVERT 1500mL TO LITERS

OpenStudy (anonymous):

Do you want the answer, or do you want to know how to do it?

OpenStudy (ghazi):

1 liter= 1000 ml 1 ml = 1/1000 L 1500 ml = 1/1000* 1500 L

OpenStudy (ghazi):

\[1500ml=\frac{ 1 }{ 1000 }*1500 L\]

OpenStudy (anonymous):

\(1500 \ ml = 1500\ mL \times \frac{1\ L}{1000\ mL}\)

OpenStudy (anonymous):

\(1500\ ml = 1500\ \cancel{mL} \times \frac{1\ L}{1000\ \cancel{ mL}}\)

OpenStudy (anonymous):

multiply together